The Fresh Connection sells cases of PET bottles to sports camps at an price of $18.81 and the materials cost an average of $5.71 per case. They make 8950 cases per week. They sell 4922 cases in a particular week, leaving them with excess that must be discarded. Now, however, you have suggested a change in the process where the camps submit an order in advance for a specific number of cases. The result will be a decrease in sales of 23 % because customers don't want to make this commitment in conjunction with a proposed price increase. You are proposing an increase in price of 14 % because they are guaranteed to receive exactly what they order each week and you want to offset some of the lost sales. An added advantage is that The Fresh Connection only has to produce exactly the number of cases that are ordered. How much better or worse off would The Fresh Connection be if they implemented your suggestion? (Unsold juice no salvage value in this scenario) Don't round any values until you reach the end of the problem and then round to two (2) decimal places.
Answer #1

Current situation

Cost of production = 8950*5.71 = 51104.5

Revenue = 4922*18.81 = 92582.82

Profit = 92582.82-51104.5 = 41478.32

New situation

Cost of production = (1-0.23)*4922*5.71 = 21640.55

Revenue = (1-0.23)*4922*(18.81*1.14) = 81269.19

Profit = 81269.19-21640.55 = 59628.64

The difference

Difference = 59628.64-41478.32 = 18150.32 better

The new strategy will yield 18150.32 higher profit